# On a nonlocal superconductivity problem

### v1

@@ 1. Introduction
Building on the framework of the aforementioned works, we formulate the problem studied in this paper. Given a smooth function $f$, we consider an unknown pair $(u,\Omega)$, where $u$ is a function defined in $\mathbb{R}^{n}$ and $\Omega\subset\mathbb{R}^{n}$ is an open set, satisfying
@@ 2. Preliminaries
Solutions are understood in the viscosity sense, for test functions $\varphi$ that touch $u$ at point $x$, satisfying $|\nu D\varphi(x)+\xi|>0$, see [21, Defintion 1.3]. Additionally, we observe that solutions for (2.2) are expected to be in $L^{1}_{s}(\mathbb{R}^{n})$, as they further satisfy inequality
@@
Lipschitz estimates for solutions of (2.2) are available, with the advantage that they are uniform with respect to $\xi\in\mathbb{S}^{n-1}$ (see [21, Lemma 2.2]). For the reader’s convenience, we present this result as follows.
@@ 3. Nonlocal positivity argument
$\Delta^{s}w_{1}(x)=f(x),\quad\mbox{for}\quad f:=-\Delta^{s}w_{2}.$
@@
$\left|\left\{x\in B_{r_{\star}}:Du(x)\cdot e\leq\delta\right\}\right|>\mu|B_{r_{\star}}|,$
@@
$\left|\left\{x\in B_{r_{\star}}:\overline{w}\geq 1-\delta\right\}\right|>\mu\left|B_{r_{\star}}\right|.$
By [11, Theorem 10.4] translated to $B_{r_{\star}}(x)$, for some $\epsilon>0$ depending on $n$ and $s$, we obtain
$\mu\,|B_{r_{\star}}|$ $<$ $\left|\left\{x\in B_{r_{\star}}:\overline{w}\geq 1-\delta\right\}\right|$
@@
for $x\in B_{r_{\star}/4}$. Rearranging terms,
$c(1-\delta)\,\mu^{\frac{1}{\epsilon}}\leq\overline{w}(x)+r_{\star}^{2s}C_{0},$
@@
$\inf_{e\in\mathbb{S}^{n-1}}\left|\left\{x\in B_{r_{\star}\lambda^{i}}:Du(x)\cdot e\leq\delta\lambda^{\alpha i}\right\}\right|>\mu|B_{r_{\star}\lambda^{i}}|,$ (3.1)
@@
$\inf_{e\in\mathbb{S}^{n-1}}\left|\left\{x\in B_{r_{\star}}:Dv_{i}(x)\cdot e\leq\delta\right\}\right|>\mu|B_{r_{\star}}|.$
@@ 4. Oscillation estimates
$|Dw(x)-Dw(0)|\leq C|x|^{2s-1}\quad\mbox{for}\quad x\in B_{1/2},$
@@
$\mathcal{A}=\left\{x\in B_{1}:Dv(x)\cdot e\leq\delta\right\}.$
@@ 5. Gradient regularity estimates
$\mathcal{A}(i)\coloneqq\inf_{e\in\mathbb{S}^{n-1}}\left|\left\{x\in B_{r_{\star}\lambda^{i}}:Dv(x)\cdot e\leq\delta\lambda^{\alpha i}\right\}\right|,$
@@
$\left|\left\{x\in B_{1}:Dw(x)\cdot e\leq\delta\right\}\right|\leq\mu|B_{1}|,$ (5.3)
@@ 6. Fully Nonlinear Operators
In this setting, following the ideas from the previous sections with appropriate adaptations for the class $\mathcal{L}_{1}(s)$, we are able to establish the following theorem.

### v2

@@ 1. Introduction
Building on the framework of the aforementioned works, we formulate the problem studied in this paper. Given a smooth function $f$, we consider an unknown pair $(u,\Omega)$, where $u$ is a function defined in $\mathbb{R}^{n}$ and $\Omega\subset\mathbb{R}^{n}$ is an open set, satisfying in the viscosity sense
@@ 1.2. Main results and consequences
The regularity assumption on $f$ is natural in view of the structure of the argument. Indeed, at a certain stage, $L^{\varepsilon}$-estimates must be applied, which are currently available only under uniform $L^{\infty}$ bounds on $\partial_{i}f$, since the PDE is differentiated in the region $B_{1}\cap\Omega$. We nevertheless believe that this condition could be weakened. However, the primary focus of the present paper is the development of sharp estimates in the homogeneous setting, which we expect to constitute a key ingredient for extending the theory to weaker regularity assumptions on $f$. Establishing such extensions is left for future investigation.
@@ 2. Preliminaries
Solutions are understood in the viscosity sense, for test functions $\varphi$ that touch $u$ at point $x$, satisfying $|\nu D\varphi(x)+\xi|>0$, see [22, Defintion 1.3]. Additionally, we observe that solutions for (2.2) are expected to be in $L^{1}_{s}(\mathbb{R}^{n})$, as they further satisfy inequality
@@
Lipschitz estimates for solutions of (2.2) are available, with the advantage that they are uniform with respect to $\xi\in\mathbb{S}^{n-1}$ (see [22, Lemma 2.2]). For the reader’s convenience, we present this result as follows.
@@
We conclude this section by recording the following $L^{\epsilon}$-estimate, which is stated to the fractional Laplacian setting in [11, Theorem 10.4] and provides a quantitative control on the measure of sublevel sets.
Lemma 2.1.
Let $r>0$ and $z\in\mathbb{R}^{n}$. Consider $u\geq 0$ in $\mathbb{R}^{n}$ and $\Delta^{s}u\leq C_{0}$ in $B_{2r}(z)$. Then, there are universal constants $C>0$ and $\epsilon>0$ such that
$|\{u<t\}\cap B_{r}(z)|\leq Cr^{n}(u(z)+C_{0}r^{2s})^{\epsilon}t^{-\epsilon}\quad\forall t>0.$
@@ 3. Nonlocal positivity argument
$\Delta^{s}w_{1}(x)=f(x),\quad\mbox{for}\quad f\coloneqq-\Delta^{s}w_{2}.$
@@
$\left|\left\{x\in B_{r_{\star}}\colon Du(x)\cdot e\leq\delta\right\}\right|>\mu|B_{r_{\star}}|,$
@@
$\left|\left\{x\in B_{r_{\star}}\colon\overline{w}\geq 1-\delta\right\}\right|>\mu\left|B_{r_{\star}}\right|.$
We now observe that for any $y\in B_{r_{\star}/4}$, the inclusion $B_{r_{\star}}\subset B_{3r_{\star}/2}(y)$ holds. Consequently, Lemma 2.1 can be applied to yield
$\mu\,|B_{r_{\star}}|$ $<$ $\left|\left\{x\in B_{r_{\star}}\colon\overline{w}\geq 1-\delta\right\}\right|$
$\leq$ $\left|\left\{x\in B_{3r_{\star}/2}(y)\colon\overline{w}\geq 1-\delta\right\}\right|$
@@
for $y\in B_{r_{\star}/4}$. Rearranging terms,
$c(1-\delta)\,\mu^{\frac{1}{\epsilon}}\leq\overline{w}(y)+r_{\star}^{2s}C_{0},$
@@
$\inf_{e\in\mathbb{S}^{n-1}}\left|\left\{x\in B_{r_{\star}\lambda^{i}}\colon Du(x)\cdot e\leq\delta\lambda^{\alpha i}\right\}\right|>\mu|B_{r_{\star}\lambda^{i}}|,$ (3.1)
@@
$\inf_{e\in\mathbb{S}^{n-1}}\left|\left\{x\in B_{r_{\star}}\colon Dv_{i}(x)\cdot e\leq\delta\right\}\right|>\mu|B_{r_{\star}}|.$
@@ 4. Oscillation estimates
$|Dw(x)-Dw(0)|\leq C|x|^{2s-1}\quad\mbox{for}\quad x\in B_{1/2},$ (4.2)
@@
$\mathcal{A}=\left\{x\in B_{1}\colon Dv(x)\cdot e\leq\delta\right\}.$
@@ 5. Gradient regularity estimates
$\mathcal{A}(i)\coloneqq\inf_{e\in\mathbb{S}^{n-1}}\left|\left\{x\in B_{r_{\star}\lambda^{i}}\colon Dv(x)\cdot e\leq\delta\lambda^{\alpha i}\right\}\right|,$
@@
$\left|\left\{x\in B_{1}\colon Dw(x)\cdot e\leq\delta\right\}\right|\leq\mu|B_{1}|,$ (5.3)
@@ 6. Fully Nonlinear Operators
The next ingredients we require are the corresponding versions of Lemma 3.2 and Proposition 4.1. Although their proofs follow the same general strategy as before, a few minor adjustments are required, and we briefly comment on them below.
First, we invoke the most general formulation of Lemma 2.1, stated in terms of the nonlocal Pucci operators (see [11, Theorem 10.4]). With this modification, the proof of Proposition 3.1 carries over unchanged.
For the proof of the analogue of Proposition 4.1, only a minor modification is required: in estimate (4.2), the exponent 2s12s-1 is replaced by a universal exponent $\overline{\alpha}\leq 2s-1$. This exponent arises from the interior regularity theory established in [12].
After these adjustments, we follow the same strategy as in the proof of Theorem 1.1, to establish the following result.